OpenStudy (anonymous):

This is such a great deal! Only $10 to join and $17.50 for each DVD, I order including shipping charges! I think I will order some DVDs. Emma joined a DVD Movie Club. It cost $10 to join and $17.50 for each DVD ordered including shipping charges. When her first order arrived, she received a bill for $115. How many DVDs did she order? A. Write an equation to solve the problem: B. Solve the problem and show all of your work:

OpenStudy (anonymous):

@iGreen @Love_Ranaa @Conqueror

OpenStudy (anonymous):

My equation -

OpenStudy (anonymous):

(17 x ?) + 10 =

OpenStudy (igreen):

$10 to join, $17.50 for each DVD

OpenStudy (igreen):

Correct..but we can replace 17 x ? with 17x.

OpenStudy (igreen):

17x + 10 = What goes after the equal sign? What was her total bill?

OpenStudy (anonymous):

Yes @green 17.50 x 6 = 105 105 + 10 = 115

OpenStudy (igreen):

No.. 17x + 10 = 115

OpenStudy (igreen):

Now solve for 'x'

OpenStudy (anonymous):

@iGreen How would I figure out "X"

OpenStudy (anonymous):

Subtract?

OpenStudy (texaschic101):

first subtract 10 from both sides of the equal sign and what do you get ?

OpenStudy (anonymous):

115 - 10 = 105 17 - 10 = 7 @texaschic101

OpenStudy (texaschic101):

17x + 10 = 115 -- subtract 10 from both sides 17x + 10 - 10 = 115 - 10 -- the 10's cancel out on the left side 17x = 105 -- now divide both sides by 17

OpenStudy (igreen):

It's 17x, not 17

OpenStudy (texaschic101):

wait...the equation is wrong

OpenStudy (anonymous):

115 - 10 = 105 17 - 10 = 7 17 divided by 7 = 2.428 107 divided by 7 = 15.28

OpenStudy (texaschic101):

17.5x + 10 = 115 is the correct equation because the CD's cost 17.50 not 17

OpenStudy (texaschic101):

17.5x + 10 = 115 -- subtract 10 from both sides 17.5x + 10 - 10 = 115 - 10 17.5x = 105 -- now divide both sides by 17.5

OpenStudy (anonymous):

17.50x + 10 = 115 Would be the right equation and my math I've done up there is how I got 6 CD'S 17.50 x 6 = 105 105 + 10 = 115 ______________________________________________________________ 17.50 - 10 = 7.5 115 - 10 = 105 17.50 / 17.50 = 1 105 / 17.50 = 6

OpenStudy (texaschic101):

yes, that is the equation...or you could just put 17.5x + 10 = 115...they are the same. And the answer you got is correct...it is 6

OpenStudy (anonymous):

Thank you! @texaschic101

OpenStudy (texaschic101):

no problem :)...sorry about the mistake earlier

OpenStudy (igreen):

Nice work! \(\Huge\color{lime}\checkmark\) \(\Huge\ddot\smile\)

OpenStudy (anonymous):

Its okay ! :D Thanks alot guys I really appreciate the help. @iGreen @texaschic101